OpenStudy (anonymous):

how do I change the limit when I do u substitution \[\int_{0}^{\frac{\pi}{2}} cos^{5}xdx\] \[\int (1-u^{2})^{2} du\]

OpenStudy (anonymous):

And which substitution are you choosing to do here? Just to be sure

OpenStudy (anonymous):

u=sinx du=cosx dx

OpenStudy (anonymous):

Alright, just checking your understanding ;-D simple: u = sin(x) so... sin(0)=0 sin(\(\frac{\pi}{2}\))=1

OpenStudy (anonymous):

If you substitute back at the end, you can change the integrand back to what it was originally. The key thing to remember is the limits of integration MUST match the variable expressed.

OpenStudy (anonymous):

thanks agent